OpenStudy (anonymous):

Find the quotient of the quantity negative 5 times x to the 3rd power plus 20 times x to the 2nd power minus 25 times x all over negative 5 times x.

OpenStudy (anonymous):

−5x^3 + 20x^2 + 5 x^2 − 4x + 5 −x^2 − 4x + 5 x^2 + 4x − 5

OpenStudy (johnweldon1993):

\[\large \frac{ -5x^3 + 20x^2 - 25x }{ -5x }\] If you can answer these 3 questions you can answer the question what is: \[\large \frac{ -5x^3 }{ 5x } = \space ?\] \[\large \frac{ 20x^2 }{ 5x } = \space ?\] \[\large \frac{ -25x }{ 5x } = \space ?\]

OpenStudy (anonymous):

hm I remember u telling me something about this

OpenStudy (anonymous):

Forgot to write that down on my notes since i dont have this one to slove for it

OpenStudy (johnweldon1993):

Right we went over something very similar... okay take it step by step \[\large \frac{ -5x^3 }{ -5x } = \frac{ -5 }{ -5 } \times \frac{ x^3 }{ x }\] remember you can just break everything up like that...so what is -5/-5? and what is x^3/x? *hint...when dividing exponents with the same base (x^3/x) you leave the base alone (x) and subtract the exponents (3 - 1)

OpenStudy (anonymous):

5^3-1? = 5^2

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (johnweldon1993):

Close but you forgot the first part -5/-5 \[\large \frac{ -5x^3 }{ -5x } = x^2\] Make sense there?